OpenStudy (anonymous):

f(x) = 8 - 2x. Find f^-1(y)

OpenStudy (anonymous):

f(x) = 8 - 2x y = 2x - 8 x = 2y - 8 x +8 = 2y f^-1(x) = x + 8 / 2 correct?

OpenStudy (anonymous):

it looks weird that I have f^-1(x) = x + 8 / 2

OpenStudy (anonymous):

@tHe_FiZiCx99 go ahead fizi :)

OpenStudy (the_fizicx99):

._. I'm watching anime ;_;

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x) Alright with that said, lets do it in steps: f(x) = 8-2x y = 8-2x x=8-2y x-8=2y (x-8)/2 = y f^-1(x) = (x-8)/2

OpenStudy (the_fizicx99):

I think you mean to say f^-1(x) And no, you went wrong in the signs. f(x) = -2x + 8 y = -2x + 8 Switch the x for the y x = -2y + 8, now solve for Y (x - 8) = -2y /-2 /-2 \(\ \sf \Large \dfrac{(x - 8)}{-2} = y \) Now change y back to\(\ \sf \Large f^{-1}(x)\)

OpenStudy (anonymous):

I switched it because 8 - 2x was different than the way I usually see it written so I changed it to 2x - 8

OpenStudy (anonymous):

I have to find f^-1(y)

OpenStudy (anonymous):

Oh the negative so yeah your final answer is f^-1(x) = (8-x)/2 oh this is why we use latex.

OpenStudy (the_fizicx99):

No ^_~ It's \(\ \sf \color{blue}{+8} ~ or ~ just~8 \) and \(\ \color{red}{-2x}\) that becomes -2x + 8 or -2x + 8

OpenStudy (anonymous):

The problem is Let f(x) = 8 - 2x. Find f^-1(y)

OpenStudy (the_fizicx99):

Hmm, never done f(y) before

OpenStudy (anonymous):

yeah, I thought it was strange

OpenStudy (the_fizicx99):

Simply because f(x) is function of x or y

OpenStudy (anonymous):

Do you think that is a typo on the worksheet?

OpenStudy (the_fizicx99):

Most likely :3 ask @Luigi0210 <-- My Older Sister knows the way ~

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

Sorry OS just stopped working, but it's the same thing just replace the x and y again.

OpenStudy (anonymous):

So instead of x's in your final answers just put y.

OpenStudy (anonymous):

I put my final answer as f^-1(x) = (x-8) / 2

OpenStudy (anonymous):

No you put f^-1(y) = (y-8)/2

OpenStudy (anonymous):

(8-y)/2 not y-8

OpenStudy (anonymous):

f^-1(y) = (8-y) / 2

OpenStudy (luigi0210):

Because he's batman~

OpenStudy (anonymous):

so it is the same steps to find the inverse function, you just replace the x's with y's?

OpenStudy (anonymous):

Yeah pretty much, just going backwards.

OpenStudy (the_fizicx99):

Erm, nope @iambatman , your answer is wrong :3

OpenStudy (anonymous):

Elaborate?

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

The 2 should be negative.

OpenStudy (anonymous):

Do you even algebra? It's the same thing.

OpenStudy (anonymous):

Oh look, mr wolfram agrees with me :) http://www.wolframalpha.com/input/?i=f%28x%29%3D8-2x+inverse